OpenStudy (anonymous):

The volume of the solid obtained by rotating the region enclosed by, x=0 y=1 and x=y^5, about the line y=1. Find the volume

OpenStudy (anonymous):

all i need is the formula. from [a,b] V=()dx

OpenStudy (anonymous):

man its been a while since i did this stuff. i got (5/6)*pi

OpenStudy (amistre64):



OpenStudy (amistre64):

the interval is from 0 to 1

OpenStudy (anonymous):

so V=pi?dx

OpenStudy (amistre64):

the radiuses are y^5

OpenStudy (amistre64):

pi [y^5]^2

OpenStudy (anonymous):

i got \[5\Pi/6\]

OpenStudy (amistre64):

\[pi\int_{0}^{1}y^{10}\ dy\] if i see it right

OpenStudy (anonymous):

i have no idea why u have to the 10th

OpenStudy (amistre64):

(y^5)^2 is the rotation

OpenStudy (anonymous):

ohh you are using y^5, not fifth root of x. cool.

OpenStudy (anonymous):

should it not be in terms of x since i have a dx?

OpenStudy (amistre64):

no, the graph integral really doesnt care what you call it; the names are just to help us position it all

OpenStudy (anonymous):

for some reason it is saying its wrong. I think it wants it in terms of x for my answer

OpenStudy (zarkon):

\[\frac{\pi}{21}\]

OpenStudy (amistre64):

i gotta see where my eyes went wrong at

OpenStudy (anonymous):

i got (10/11)*pi ..

OpenStudy (amistre64):

the radius function in terms of x will be: 1 - x^(1/5) right?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

\[\int\limits_{0}^{1}\pi(1-x^{1/5})^2dx\] or \[\int\limits_{0}^{1}2\pi y^5(1-y)dy\]

OpenStudy (amistre64):

1 - x^(1/5) 1 - x^(1/5) ---------- 1 - x^(1/5) -x^(1/5) + x^(2/5) -------------------- 1 -2x^{1/5} + x^{2/5} \[pi\int_{0}^{1} ( 1 -2x^{1/5} + x^{2/5})dx\] \[pi (x -\frac{5x^{6/5}}{3} + \frac{5x^{7/5}}{7})\] \[pi (1 -\frac{5}{3} + \frac{5}{7})\]

OpenStudy (amistre64):

maybe?

OpenStudy (anonymous):

we've all been out of the pre-calc game for too long!!

OpenStudy (amistre64):

my first thought was mixing shells and discs lol